OpenStudy (pottersheep):

Trig Identities....Half Angles help PLEASEEEEE!!! Can someone explain what the teacher wrote on the board for me please??? Cos2A = cos^2A-sin^2A = 1 - 2sin^2 A (What happened here???????)

OpenStudy (anonymous):

sin^2 + cos^2 =1

OpenStudy (anonymous):

1-cos^2==sin^2

OpenStudy (anonymous):

sub in ... Cos2A=cos^2A-1+cos^2A

OpenStudy (anonymous):

Cos2A=2cos^2A -1

OpenStudy (anonymous):

now do same thing with 1-sin^2= cos^2 and you should get the identity you teacher wrote sooooo 1-sin^2=cos^2 sub into Cos2A = cos^2A-sin^2A Cos2A=1-sin^2A-sin^A ==1-2sin^2A

OpenStudy (pottersheep):

HMmm let me process what you said

OpenStudy (pottersheep):

Thank you! I finally get it!

OpenStudy (anonymous):

yw